I have two questions to ask.
1. what annual rate of interest is required to triple an investment in 8 years?
2.Solve 4(7-3x)=(1/16)

Answer by stanbon(75887)

1. what annual rate of interest is required to triple an investment in 8 years?
Question: Under what kind of compounding?
----------------------------------------------
2. Solve
4(7-3x)=(1/16)
(7-3x) = 1/64
2x = 7-1/64
---
2x = (7*64-1)/64
---
2x = 447/64
---
x = 447/128
---
x = 3.4922
================
Cheers,
Stan H.
================